The equation y = -1777x + 27,153 can be used to predict the number y of gun deaths in the United States x years after 2000, that is, x = 0 corresponds to 2000, x = 2 corresponds to 2002, x = 5 corresponds to 2005, and so on. Predict the number of gun deaths in 2005 and 2007. In what year will the number of gun deaths be 12,937?
The predicted gun deaths in 2005 will be?
2007?
The predicted number of gun deaths will be 12,937 will be in what year?

Answer by oscargut(2103)

12,937=-1777x+27,153 then 1777x=27153-12937=14216
then x=14216/1777=8
Answer: in 2008 the number of gun deaths will be 12,937
-1777(5)+27153=18268
the predicted number of gun death in 2005 will be =18268
-1777(7)+27153=14714
the predicted number of gun death in 2007 will be =14714